OpenStudy (vortish):

solve the equation for y to obtain the slope intecept form of the equation 3x + y = 1 x -y =7

OpenStudy (ny,ny):

to solve for y, you have to get y on one side of the equal sign. so... 3x + y = 1 y = -3x +1 and x - y = 7 -y = -x + 7 divide by -1 y = x - 7

OpenStudy (ny,ny):

understand? :o

OpenStudy (vortish):

yes

OpenStudy (ny,ny):

ok good. :)